# Context is Key: A Benchmark for Forecasting with Essential Textual Information

### v3

@@ C.3 Full results partitioned by model capabilities
    Llama-3.1-405B-Inst 0.159 $\pm$ 0.008 7.440 $\pm$ 0.586 0.140 $\pm$ 0.013 0.109 $\pm$ 0.002 0.191 $\pm$ 0.006 0.133 $\pm$ 0.001 0.167 $\pm$ 0.008 0.316 $\pm$ 0.028 0.376 $\pm$ 0.039
    Llama-3-70B-Inst 0.286 $\pm$ 0.004 14.081 $\pm$ 0.195 0.323 $\pm$ 0.008 0.122 $\pm$ 0.003 0.408 $\pm$ 0.012 0.168 $\pm$ 0.002 0.492 $\pm$ 0.019 0.473 $\pm$ 0.012 0.577 $\pm$ 0.017
    Llama-3-8B-Inst 0.461 $\pm$ 0.008 26.913 $\pm$ 0.559 0.323 $\pm$ 0.010 0.174 $\pm$ 0.003 0.849 $\pm$ 0.021 0.406 $\pm$ 0.014 1.245 $\pm$ 0.039 0.437 $\pm$ 0.004 0.494 $\pm$ 0.004
    Mixtral-8x7B-Inst 0.523 $\pm$ 0.023 31.389 $\pm$ 0.388 0.433 $\pm$ 0.043 0.204 $\pm$ 0.000 0.864 $\pm$ 0.029 0.427 $\pm$ 0.024 1.245 $\pm$ 0.006 0.644 $\pm$ 0.080 0.789 $\pm$ 0.112
    Qwen-2.5-7B-Inst 0.289 $\pm$ 0.003 21.816 $\pm$ 0.724 0.343 $\pm$ 0.005 0.127 $\pm$ 0.002 0.324 $\pm$ 0.008 0.204 $\pm$ 0.005 0.281 $\pm$ 0.014 0.409 $\pm$ 0.002 0.480 $\pm$ 0.002
    Qwen-2.5-1.5B-Inst 0.575 $\pm$ 0.014 30.636 $\pm$ 0.834 0.610 $\pm$ 0.007 0.214 $\pm$ 0.004 0.988 $\pm$ 0.049 0.341 $\pm$ 0.009 1.077 $\pm$ 0.122 0.896 $\pm$ 0.006 1.151 $\pm$ 0.008
    Qwen-2.5-0.5B-Inst 0.464 $\pm$ 0.012 26.261 $\pm$ 0.513 0.268 $\pm$ 0.008 1.029 $\pm$ 0.055 0.744 $\pm$ 0.039 0.244 $\pm$ 0.007 2.043 $\pm$ 0.104 0.330 $\pm$ 0.004 0.392 $\pm$ 0.005
    GPT-4o 0.276 $\pm$ 0.010 8.601 $\pm$ 0.410 0.180 $\pm$ 0.004 0.087 $\pm$ 0.003 0.519 $\pm$ 0.029 0.113 $\pm$ 0.006 0.447 $\pm$ 0.029 0.590 $\pm$ 0.033 0.769 $\pm$ 0.046
    GPT-4o-mini 0.353 $\pm$ 0.022 16.461 $\pm$ 0.462 0.296 $\pm$ 0.043 0.419 $\pm$ 0.014 0.471 $\pm$ 0.012 0.218 $\pm$ 0.005 1.024 $\pm$ 0.033 0.475 $\pm$ 0.080 0.578 $\pm$ 0.112
@@
    Llama-3-70B-Inst 0.550 $\pm$ 0.013 17.043 $\pm$ 0.535 0.645 $\pm$ 0.018 0.284 $\pm$ 0.015 0.392 $\pm$ 0.014 0.519 $\pm$ 0.026 0.312 $\pm$ 0.019 0.453 $\pm$ 0.020 0.495 $\pm$ 0.028
    Llama-3-70B 0.237 $\pm$ 0.006 11.960 $\pm$ 0.668 0.310 $\pm$ 0.011 0.126 $\pm$ 0.009 0.217 $\pm$ 0.007 0.134 $\pm$ 0.003 0.241 $\pm$ 0.019 0.294 $\pm$ 0.008 0.329 $\pm$ 0.010
    Llama-3-8B-Inst 0.484 $\pm$ 0.010 18.216 $\pm$ 0.443 0.345 $\pm$ 0.002 0.138 $\pm$ 0.004 0.910 $\pm$ 0.030 0.242 $\pm$ 0.008 1.278 $\pm$ 0.069 0.617 $\pm$ 0.022 0.787 $\pm$ 0.030
    Llama-3-8B 0.313 $\pm$ 0.023 18.358 $\pm$ 0.947 0.404 $\pm$ 0.043 0.124 $\pm$ 0.003 0.280 $\pm$ 0.026 0.179 $\pm$ 0.014 0.267 $\pm$ 0.015 0.530 $\pm$ 0.084 0.661 $\pm$ 0.117
    Mixtral-8x7B-Inst 0.264 $\pm$ 0.004 15.210 $\pm$ 0.605 0.344 $\pm$ 0.004 0.127 $\pm$ 0.003 0.224 $\pm$ 0.005 0.179 $\pm$ 0.010 0.173 $\pm$ 0.009 0.348 $\pm$ 0.005 0.405 $\pm$ 0.007
    Mixtral-8x7B 0.262 $\pm$ 0.008 15.890 $\pm$ 0.481 0.348 $\pm$ 0.012 0.146 $\pm$ 0.022 0.230 $\pm$ 0.016 0.153 $\pm$ 0.002 0.230 $\pm$ 0.041 0.354 $\pm$ 0.007 0.414 $\pm$ 0.009
    Qwen-2.5-7B-Inst 1.980 $\pm$ 0.027 42.505 $\pm$ 0.715 1.825 $\pm$ 0.048 1.022 $\pm$ 0.054 2.215 $\pm$ 0.046 2.761 $\pm$ 0.024 1.723 $\pm$ 0.092 2.025 $\pm$ 0.093 1.607 $\pm$ 0.127
    Qwen-2.5-7B 0.904 $\pm$ 0.037 35.488 $\pm$ 0.975 0.690 $\pm$ 0.063 0.794 $\pm$ 0.083 1.558 $\pm$ 0.062 0.879 $\pm$ 0.028 2.328 $\pm$ 0.153 0.878 $\pm$ 0.084 0.881 $\pm$ 0.113
    Qwen-2.5-1.5B-Inst 2.153 $\pm$ 0.027 47.555 $\pm$ 0.835 2.052 $\pm$ 0.046 1.566 $\pm$ 0.033 2.671 $\pm$ 0.038 2.156 $\pm$ 0.035 3.635 $\pm$ 0.053 2.480 $\pm$ 0.085 2.323 $\pm$ 0.113
    Qwen-2.5-1.5B 1.731 $\pm$ 0.036 42.324 $\pm$ 0.515 1.343 $\pm$ 0.061 1.737 $\pm$ 0.074 2.594 $\pm$ 0.042 2.256 $\pm$ 0.042 3.275 $\pm$ 0.132 2.036 $\pm$ 0.083 1.526 $\pm$ 0.114
    Qwen-2.5-0.5B-Inst 1.938 $\pm$ 0.024 47.508 $\pm$ 0.586 1.743 $\pm$ 0.043 1.800 $\pm$ 0.021 2.193 $\pm$ 0.025 2.303 $\pm$ 0.028 3.439 $\pm$ 0.004 1.685 $\pm$ 0.084 1.398 $\pm$ 0.114
    Qwen-2.5-0.5B 1.991 $\pm$ 0.024 45.726 $\pm$ 0.797 1.827 $\pm$ 0.045 0.950 $\pm$ 0.025 1.967 $\pm$ 0.020 2.799 $\pm$ 0.022 1.804 $\pm$ 0.036 1.695 $\pm$ 0.085 1.443 $\pm$ 0.113
@@
    UniTime 0.370 $\pm$ 0.001 33.377 $\pm$ 0.151 0.267 $\pm$ 0.002 0.179 $\pm$ 0.001 0.321 $\pm$ 0.001 0.511 $\pm$ 0.003 0.337 $\pm$ 0.001 0.333 $\pm$ 0.001 0.385 $\pm$ 0.001
    Time-LLM (ETTh1) 0.476 $\pm$ 0.001 41.641 $\pm$ 0.142 0.448 $\pm$ 0.002 0.192 $\pm$ 0.000 0.373 $\pm$ 0.000 0.538 $\pm$ 0.003 0.397 $\pm$ 0.001 0.382 $\pm$ 0.001 0.440 $\pm$ 0.001
@@
    Llama-3.1-405B-Inst 0.473 $\pm$ 0.005 32.450 $\pm$ 0.278 0.393 $\pm$ 0.002 0.325 $\pm$ 0.006 0.752 $\pm$ 0.015 0.494 $\pm$ 0.009 0.720 $\pm$ 0.027 0.594 $\pm$ 0.006 0.617 $\pm$ 0.006
    Llama-3-70B-Inst 0.470 $\pm$ 0.008 33.486 $\pm$ 0.348 0.404 $\pm$ 0.002 0.304 $\pm$ 0.007 0.717 $\pm$ 0.015 0.488 $\pm$ 0.022 0.694 $\pm$ 0.024 0.606 $\pm$ 0.007 0.631 $\pm$ 0.007
    Llama-3-8B-Inst 0.601 $\pm$ 0.006 39.657 $\pm$ 0.383 0.391 $\pm$ 0.005 0.322 $\pm$ 0.004 1.123 $\pm$ 0.018 0.642 $\pm$ 0.012 1.446 $\pm$ 0.035 0.581 $\pm$ 0.005 0.617 $\pm$ 0.005
    Mixtral-8x7B-Inst 0.712 $\pm$ 0.021 44.240 $\pm$ 0.357 0.625 $\pm$ 0.043 0.324 $\pm$ 0.006 1.053 $\pm$ 0.007 0.783 $\pm$ 0.005 1.237 $\pm$ 0.007 0.858 $\pm$ 0.080 0.872 $\pm$ 0.112
    Qwen-2.5-7B-Inst 0.420 $\pm$ 0.022 35.132 $\pm$ 0.508 0.447 $\pm$ 0.043 0.259 $\pm$ 0.008 0.505 $\pm$ 0.010 0.374 $\pm$ 0.007 0.411 $\pm$ 0.016 0.692 $\pm$ 0.080 0.792 $\pm$ 0.112
    Qwen-2.5-1.5B-Inst 0.450 $\pm$ 0.007 34.470 $\pm$ 0.592 0.377 $\pm$ 0.006 0.232 $\pm$ 0.004 0.661 $\pm$ 0.018 0.387 $\pm$ 0.013 0.939 $\pm$ 0.039 0.423 $\pm$ 0.007 0.476 $\pm$ 0.008
    Qwen-2.5-0.5B-Inst 0.305 $\pm$ 0.007 25.473 $\pm$ 0.325 0.267 $\pm$ 0.010 0.162 $\pm$ 0.001 0.384 $\pm$ 0.008 0.300 $\pm$ 0.011 0.440 $\pm$ 0.016 0.315 $\pm$ 0.002 0.367 $\pm$ 0.002
    GPT-4o 0.441 $\pm$ 0.008 29.858 $\pm$ 0.359 0.381 $\pm$ 0.002 0.179 $\pm$ 0.002 0.692 $\pm$ 0.028 0.356 $\pm$ 0.007 0.953 $\pm$ 0.067 0.422 $\pm$ 0.003 0.471 $\pm$ 0.002
    GPT-4o-mini 0.423 $\pm$ 0.006 34.039 $\pm$ 0.264 0.360 $\pm$ 0.003 0.214 $\pm$ 0.002 0.649 $\pm$ 0.019 0.392 $\pm$ 0.008 0.771 $\pm$ 0.041 0.461 $\pm$ 0.003 0.511 $\pm$ 0.003
@@
    Llama-3-70B-Inst 0.378 $\pm$ 0.004 24.665 $\pm$ 0.445 0.369 $\pm$ 0.003 0.150 $\pm$ 0.003 0.513 $\pm$ 0.012 0.292 $\pm$ 0.006 0.668 $\pm$ 0.025 0.384 $\pm$ 0.003 0.440 $\pm$ 0.003
    Llama-3-70B 0.312 $\pm$ 0.006 20.965 $\pm$ 0.435 0.349 $\pm$ 0.011 0.141 $\pm$ 0.002 0.351 $\pm$ 0.008 0.216 $\pm$ 0.004 0.395 $\pm$ 0.017 0.372 $\pm$ 0.004 0.434 $\pm$ 0.005
    Llama-3-8B-Inst 0.503 $\pm$ 0.009 29.628 $\pm$ 0.420 0.385 $\pm$ 0.003 0.159 $\pm$ 0.002 0.914 $\pm$ 0.030 0.430 $\pm$ 0.010 1.271 $\pm$ 0.070 0.424 $\pm$ 0.003 0.467 $\pm$ 0.001
    Llama-3-8B 0.345 $\pm$ 0.003 23.990 $\pm$ 0.356 0.326 $\pm$ 0.004 0.150 $\pm$ 0.002 0.497 $\pm$ 0.009 0.267 $\pm$ 0.005 0.640 $\pm$ 0.017 0.381 $\pm$ 0.002 0.444 $\pm$ 0.002
    Mixtral-8x7B-Inst 0.383 $\pm$ 0.015 23.531 $\pm$ 0.439 0.357 $\pm$ 0.003 0.550 $\pm$ 0.072 0.459 $\pm$ 0.047 0.211 $\pm$ 0.008 1.027 $\pm$ 0.133 0.371 $\pm$ 0.004 0.428 $\pm$ 0.004
    Mixtral-8x7B 0.306 $\pm$ 0.007 21.673 $\pm$ 0.469 0.360 $\pm$ 0.013 0.146 $\pm$ 0.002 0.327 $\pm$ 0.008 0.202 $\pm$ 0.005 0.340 $\pm$ 0.016 0.382 $\pm$ 0.002 0.445 $\pm$ 0.003
    Qwen-2.5-7B-Inst 1.023 $\pm$ 0.026 37.612 $\pm$ 1.177 0.520 $\pm$ 0.044 1.157 $\pm$ 0.022 1.634 $\pm$ 0.017 1.072 $\pm$ 0.035 3.319 $\pm$ 0.037 0.853 $\pm$ 0.081 0.769 $\pm$ 0.112
    Qwen-2.5-7B 0.732 $\pm$ 0.030 35.801 $\pm$ 1.101 0.649 $\pm$ 0.048 0.359 $\pm$ 0.036 0.974 $\pm$ 0.037 0.752 $\pm$ 0.039 1.433 $\pm$ 0.096 0.728 $\pm$ 0.084 0.730 $\pm$ 0.112
    Qwen-2.5-1.5B-Inst 1.515 $\pm$ 0.033 44.209 $\pm$ 1.056 1.314 $\pm$ 0.055 1.159 $\pm$ 0.057 1.802 $\pm$ 0.031 1.653 $\pm$ 0.042 3.383 $\pm$ 0.106 1.108 $\pm$ 0.082 0.848 $\pm$ 0.112
    Qwen-2.5-1.5B 1.069 $\pm$ 0.028 38.510 $\pm$ 1.025 1.005 $\pm$ 0.048 0.287 $\pm$ 0.026 1.339 $\pm$ 0.023 1.264 $\pm$ 0.040 1.798 $\pm$ 0.047 0.771 $\pm$ 0.086 0.720 $\pm$ 0.112
    Qwen-2.5-0.5B-Inst 1.318 $\pm$ 0.037 41.697 $\pm$ 0.715 1.463 $\pm$ 0.064 0.239 $\pm$ 0.031 1.192 $\pm$ 0.019 1.434 $\pm$ 0.047 1.675 $\pm$ 0.072 0.930 $\pm$ 0.082 0.743 $\pm$ 0.112
    Qwen-2.5-0.5B 1.819 $\pm$ 0.027 45.091 $\pm$ 0.689 1.707 $\pm$ 0.045 0.572 $\pm$ 0.040 1.722 $\pm$ 0.033 2.492 $\pm$ 0.036 1.671 $\pm$ 0.064 1.492 $\pm$ 0.083 1.113 $\pm$ 0.114
@@
    UniTime 0.405 $\pm$ 0.002 35.046 $\pm$ 0.176 0.361 $\pm$ 0.002 0.166 $\pm$ 0.001 0.319 $\pm$ 0.001 0.496 $\pm$ 0.005 0.314 $\pm$ 0.002 0.352 $\pm$ 0.002 0.409 $\pm$ 0.003
    Time-LLM (ETTh1) 0.458 $\pm$ 0.002 40.647 $\pm$ 0.155 0.440 $\pm$ 0.003 0.191 $\pm$ 0.000 0.371 $\pm$ 0.000 0.499 $\pm$ 0.002 0.399 $\pm$ 0.001 0.368 $\pm$ 0.001 0.424 $\pm$ 0.001

### v4

@@ C.3 Full results partitioned by model capabilities
    Llama-3.1-405B-Inst 0.159 $\pm$ 0.008 7.971 $\pm$ 0.585 0.140 $\pm$ 0.013 0.109 $\pm$ 0.002 0.191 $\pm$ 0.006 0.132 $\pm$ 0.001 0.167 $\pm$ 0.008 0.316 $\pm$ 0.028 0.376 $\pm$ 0.039
    Llama-3-70B-Inst 0.286 $\pm$ 0.004 14.802 $\pm$ 0.203 0.323 $\pm$ 0.008 0.122 $\pm$ 0.003 0.408 $\pm$ 0.012 0.168 $\pm$ 0.002 0.492 $\pm$ 0.019 0.473 $\pm$ 0.012 0.577 $\pm$ 0.017
    Llama-3-8B-Inst 0.461 $\pm$ 0.008 28.016 $\pm$ 0.584 0.323 $\pm$ 0.010 0.174 $\pm$ 0.003 0.849 $\pm$ 0.021 0.407 $\pm$ 0.014 1.245 $\pm$ 0.039 0.437 $\pm$ 0.004 0.494 $\pm$ 0.004
    Mixtral-8x7B-Inst 0.523 $\pm$ 0.023 33.069 $\pm$ 0.413 0.433 $\pm$ 0.043 0.204 $\pm$ 0.000 0.864 $\pm$ 0.029 0.426 $\pm$ 0.024 1.245 $\pm$ 0.006 0.644 $\pm$ 0.080 0.789 $\pm$ 0.112
    Qwen-2.5-7B-Inst 0.290 $\pm$ 0.003 22.852 $\pm$ 0.734 0.343 $\pm$ 0.005 0.127 $\pm$ 0.002 0.324 $\pm$ 0.008 0.205 $\pm$ 0.005 0.281 $\pm$ 0.014 0.409 $\pm$ 0.002 0.480 $\pm$ 0.002
    Qwen-2.5-1.5B-Inst 0.575 $\pm$ 0.014 32.631 $\pm$ 0.861 0.610 $\pm$ 0.007 0.214 $\pm$ 0.004 0.988 $\pm$ 0.049 0.344 $\pm$ 0.009 1.077 $\pm$ 0.122 0.896 $\pm$ 0.006 1.151 $\pm$ 0.008
    Qwen-2.5-0.5B-Inst 0.463 $\pm$ 0.012 27.221 $\pm$ 0.533 0.267 $\pm$ 0.008 1.029 $\pm$ 0.055 0.744 $\pm$ 0.039 0.244 $\pm$ 0.007 2.043 $\pm$ 0.104 0.330 $\pm$ 0.004 0.392 $\pm$ 0.005
    GPT-4o 0.274 $\pm$ 0.010 8.640 $\pm$ 0.436 0.179 $\pm$ 0.004 0.087 $\pm$ 0.003 0.519 $\pm$ 0.029 0.110 $\pm$ 0.006 0.447 $\pm$ 0.029 0.590 $\pm$ 0.033 0.769 $\pm$ 0.046
    GPT-4o-mini 0.354 $\pm$ 0.022 17.573 $\pm$ 0.505 0.296 $\pm$ 0.043 0.419 $\pm$ 0.014 0.471 $\pm$ 0.012 0.219 $\pm$ 0.005 1.024 $\pm$ 0.033 0.475 $\pm$ 0.080 0.578 $\pm$ 0.112
@@
    Llama-3-70B-Inst 0.539 $\pm$ 0.013 18.042 $\pm$ 0.572 0.641 $\pm$ 0.018 0.284 $\pm$ 0.015 0.392 $\pm$ 0.014 0.495 $\pm$ 0.025 0.312 $\pm$ 0.019 0.453 $\pm$ 0.020 0.495 $\pm$ 0.028
    Llama-3-70B 0.236 $\pm$ 0.006 12.377 $\pm$ 0.723 0.309 $\pm$ 0.011 0.126 $\pm$ 0.009 0.217 $\pm$ 0.007 0.132 $\pm$ 0.003 0.241 $\pm$ 0.019 0.294 $\pm$ 0.008 0.329 $\pm$ 0.010
    Llama-3-8B-Inst 0.483 $\pm$ 0.010 18.585 $\pm$ 0.477 0.345 $\pm$ 0.002 0.138 $\pm$ 0.004 0.910 $\pm$ 0.030 0.242 $\pm$ 0.008 1.278 $\pm$ 0.069 0.617 $\pm$ 0.022 0.787 $\pm$ 0.030
    Llama-3-8B 0.311 $\pm$ 0.023 18.634 $\pm$ 1.039 0.403 $\pm$ 0.043 0.124 $\pm$ 0.003 0.280 $\pm$ 0.026 0.177 $\pm$ 0.014 0.267 $\pm$ 0.015 0.530 $\pm$ 0.084 0.661 $\pm$ 0.117
    Mixtral-8x7B-Inst 0.264 $\pm$ 0.004 16.078 $\pm$ 0.666 0.344 $\pm$ 0.004 0.127 $\pm$ 0.003 0.224 $\pm$ 0.005 0.179 $\pm$ 0.010 0.173 $\pm$ 0.009 0.348 $\pm$ 0.005 0.405 $\pm$ 0.007
    Mixtral-8x7B 0.262 $\pm$ 0.008 16.302 $\pm$ 0.523 0.348 $\pm$ 0.012 0.146 $\pm$ 0.022 0.230 $\pm$ 0.016 0.153 $\pm$ 0.002 0.230 $\pm$ 0.041 0.354 $\pm$ 0.007 0.414 $\pm$ 0.009
    Qwen-2.5-7B-Inst 1.974 $\pm$ 0.027 45.233 $\pm$ 0.739 1.816 $\pm$ 0.048 1.022 $\pm$ 0.054 2.215 $\pm$ 0.046 2.758 $\pm$ 0.024 1.723 $\pm$ 0.092 2.025 $\pm$ 0.093 1.607 $\pm$ 0.127
    Qwen-2.5-7B 0.910 $\pm$ 0.037 38.157 $\pm$ 1.041 0.691 $\pm$ 0.063 0.794 $\pm$ 0.083 1.558 $\pm$ 0.062 0.893 $\pm$ 0.028 2.328 $\pm$ 0.153 0.878 $\pm$ 0.084 0.881 $\pm$ 0.113
    Qwen-2.5-1.5B-Inst 2.158 $\pm$ 0.027 50.654 $\pm$ 0.863 2.056 $\pm$ 0.046 1.566 $\pm$ 0.033 2.671 $\pm$ 0.038 2.165 $\pm$ 0.035 3.635 $\pm$ 0.053 2.480 $\pm$ 0.085 2.323 $\pm$ 0.113
    Qwen-2.5-1.5B 1.731 $\pm$ 0.036 45.118 $\pm$ 0.528 1.343 $\pm$ 0.061 1.737 $\pm$ 0.074 2.594 $\pm$ 0.042 2.256 $\pm$ 0.042 3.275 $\pm$ 0.132 2.036 $\pm$ 0.083 1.526 $\pm$ 0.114
    Qwen-2.5-0.5B-Inst 1.937 $\pm$ 0.024 50.482 $\pm$ 0.612 1.740 $\pm$ 0.043 1.800 $\pm$ 0.021 2.193 $\pm$ 0.025 2.305 $\pm$ 0.028 3.439 $\pm$ 0.004 1.685 $\pm$ 0.084 1.398 $\pm$ 0.114
    Qwen-2.5-0.5B 1.995 $\pm$ 0.024 48.507 $\pm$ 0.840 1.829 $\pm$ 0.045 0.950 $\pm$ 0.025 1.967 $\pm$ 0.020 2.809 $\pm$ 0.023 1.804 $\pm$ 0.036 1.695 $\pm$ 0.085 1.443 $\pm$ 0.113
@@
    UniTime 0.370 $\pm$ 0.001 35.453 $\pm$ 0.152 0.267 $\pm$ 0.002 0.179 $\pm$ 0.001 0.321 $\pm$ 0.001 0.511 $\pm$ 0.003 0.337 $\pm$ 0.001 0.333 $\pm$ 0.001 0.385 $\pm$ 0.001
    Time-LLM (ETTh1) 0.476 $\pm$ 0.001 44.084 $\pm$ 0.150 0.448 $\pm$ 0.002 0.192 $\pm$ 0.000 0.373 $\pm$ 0.000 0.538 $\pm$ 0.003 0.397 $\pm$ 0.001 0.382 $\pm$ 0.001 0.440 $\pm$ 0.001
    ChatTime-Base 0.735 $\pm$ 0.002 39.033 $\pm$ 0.312 0.843 $\pm$ 0.002 0.216 $\pm$ 0.002 1.099 $\pm$ 0.002 0.263 $\pm$ 0.004 0.374 $\pm$ 0.004 1.788 $\pm$ 0.002 2.407 $\pm$ 0.002
    ChatTime-Chat 0.747 $\pm$ 0.005 34.182 $\pm$ 0.389 0.825 $\pm$ 0.004 0.299 $\pm$ 0.020 1.198 $\pm$ 0.015 0.305 $\pm$ 0.012 0.277 $\pm$ 0.003 2.015 $\pm$ 0.016 2.691 $\pm$ 0.022
@@
    Llama-3.1-405B-Inst 0.473 $\pm$ 0.005 34.336 $\pm$ 0.294 0.393 $\pm$ 0.002 0.325 $\pm$ 0.006 0.752 $\pm$ 0.015 0.494 $\pm$ 0.009 0.720 $\pm$ 0.027 0.594 $\pm$ 0.006 0.617 $\pm$ 0.006
    Llama-3-70B-Inst 0.470 $\pm$ 0.008 35.143 $\pm$ 0.357 0.404 $\pm$ 0.002 0.304 $\pm$ 0.007 0.717 $\pm$ 0.015 0.488 $\pm$ 0.022 0.694 $\pm$ 0.024 0.606 $\pm$ 0.007 0.631 $\pm$ 0.007
    Llama-3-8B-Inst 0.602 $\pm$ 0.006 41.873 $\pm$ 0.398 0.390 $\pm$ 0.005 0.322 $\pm$ 0.004 1.123 $\pm$ 0.018 0.643 $\pm$ 0.012 1.446 $\pm$ 0.035 0.581 $\pm$ 0.005 0.617 $\pm$ 0.005
    Mixtral-8x7B-Inst 0.712 $\pm$ 0.021 46.809 $\pm$ 0.376 0.624 $\pm$ 0.043 0.324 $\pm$ 0.006 1.053 $\pm$ 0.007 0.783 $\pm$ 0.005 1.237 $\pm$ 0.007 0.858 $\pm$ 0.080 0.872 $\pm$ 0.112
    Qwen-2.5-7B-Inst 0.421 $\pm$ 0.022 37.154 $\pm$ 0.546 0.447 $\pm$ 0.043 0.259 $\pm$ 0.008 0.505 $\pm$ 0.010 0.375 $\pm$ 0.007 0.411 $\pm$ 0.016 0.692 $\pm$ 0.080 0.792 $\pm$ 0.112
    Qwen-2.5-1.5B-Inst 0.450 $\pm$ 0.007 36.308 $\pm$ 0.619 0.377 $\pm$ 0.006 0.232 $\pm$ 0.004 0.661 $\pm$ 0.018 0.387 $\pm$ 0.013 0.939 $\pm$ 0.039 0.423 $\pm$ 0.007 0.476 $\pm$ 0.008
    Qwen-2.5-0.5B-Inst 0.305 $\pm$ 0.007 26.819 $\pm$ 0.351 0.267 $\pm$ 0.010 0.162 $\pm$ 0.001 0.384 $\pm$ 0.008 0.300 $\pm$ 0.011 0.440 $\pm$ 0.016 0.315 $\pm$ 0.002 0.367 $\pm$ 0.002
    GPT-4o 0.441 $\pm$ 0.008 31.507 $\pm$ 0.385 0.381 $\pm$ 0.002 0.179 $\pm$ 0.002 0.692 $\pm$ 0.028 0.357 $\pm$ 0.007 0.953 $\pm$ 0.067 0.422 $\pm$ 0.003 0.471 $\pm$ 0.002
    GPT-4o-mini 0.423 $\pm$ 0.006 35.715 $\pm$ 0.275 0.359 $\pm$ 0.003 0.214 $\pm$ 0.002 0.649 $\pm$ 0.019 0.391 $\pm$ 0.008 0.771 $\pm$ 0.041 0.461 $\pm$ 0.003 0.511 $\pm$ 0.003
@@
    Llama-3-70B-Inst 0.378 $\pm$ 0.004 26.031 $\pm$ 0.474 0.368 $\pm$ 0.003 0.150 $\pm$ 0.003 0.513 $\pm$ 0.012 0.292 $\pm$ 0.006 0.668 $\pm$ 0.025 0.384 $\pm$ 0.003 0.440 $\pm$ 0.003
    Llama-3-70B 0.311 $\pm$ 0.006 21.810 $\pm$ 0.464 0.349 $\pm$ 0.011 0.141 $\pm$ 0.002 0.351 $\pm$ 0.008 0.215 $\pm$ 0.004 0.395 $\pm$ 0.017 0.372 $\pm$ 0.004 0.434 $\pm$ 0.005
    Llama-3-8B-Inst 0.503 $\pm$ 0.009 31.156 $\pm$ 0.438 0.385 $\pm$ 0.003 0.159 $\pm$ 0.002 0.914 $\pm$ 0.030 0.431 $\pm$ 0.010 1.271 $\pm$ 0.070 0.424 $\pm$ 0.003 0.467 $\pm$ 0.001
    Llama-3-8B 0.345 $\pm$ 0.003 25.067 $\pm$ 0.381 0.326 $\pm$ 0.004 0.150 $\pm$ 0.002 0.497 $\pm$ 0.009 0.266 $\pm$ 0.005 0.640 $\pm$ 0.017 0.381 $\pm$ 0.002 0.444 $\pm$ 0.002
    Mixtral-8x7B-Inst 0.383 $\pm$ 0.015 24.582 $\pm$ 0.471 0.357 $\pm$ 0.003 0.550 $\pm$ 0.072 0.459 $\pm$ 0.047 0.211 $\pm$ 0.008 1.027 $\pm$ 0.133 0.371 $\pm$ 0.004 0.428 $\pm$ 0.004
    Mixtral-8x7B 0.306 $\pm$ 0.007 22.560 $\pm$ 0.512 0.360 $\pm$ 0.013 0.146 $\pm$ 0.002 0.327 $\pm$ 0.008 0.202 $\pm$ 0.005 0.340 $\pm$ 0.016 0.382 $\pm$ 0.002 0.445 $\pm$ 0.003
    Qwen-2.5-7B-Inst 1.020 $\pm$ 0.026 39.942 $\pm$ 1.233 0.521 $\pm$ 0.045 1.157 $\pm$ 0.022 1.634 $\pm$ 0.017 1.061 $\pm$ 0.034 3.319 $\pm$ 0.037 0.853 $\pm$ 0.081 0.769 $\pm$ 0.112
    Qwen-2.5-7B 0.732 $\pm$ 0.030 38.109 $\pm$ 1.185 0.649 $\pm$ 0.048 0.359 $\pm$ 0.036 0.974 $\pm$ 0.037 0.751 $\pm$ 0.039 1.433 $\pm$ 0.096 0.728 $\pm$ 0.084 0.730 $\pm$ 0.112
    Qwen-2.5-1.5B-Inst 1.515 $\pm$ 0.033 47.071 $\pm$ 1.066 1.316 $\pm$ 0.055 1.159 $\pm$ 0.057 1.802 $\pm$ 0.031 1.652 $\pm$ 0.042 3.383 $\pm$ 0.106 1.108 $\pm$ 0.082 0.848 $\pm$ 0.112
    Qwen-2.5-1.5B 1.070 $\pm$ 0.028 41.187 $\pm$ 1.053 1.005 $\pm$ 0.048 0.287 $\pm$ 0.026 1.339 $\pm$ 0.023 1.264 $\pm$ 0.040 1.798 $\pm$ 0.047 0.771 $\pm$ 0.086 0.720 $\pm$ 0.112
    Qwen-2.5-0.5B-Inst 1.318 $\pm$ 0.037 44.393 $\pm$ 0.750 1.464 $\pm$ 0.064 0.239 $\pm$ 0.031 1.192 $\pm$ 0.019 1.433 $\pm$ 0.047 1.675 $\pm$ 0.072 0.930 $\pm$ 0.082 0.743 $\pm$ 0.112
    Qwen-2.5-0.5B 1.821 $\pm$ 0.027 47.763 $\pm$ 0.715 1.705 $\pm$ 0.045 0.572 $\pm$ 0.040 1.722 $\pm$ 0.033 2.498 $\pm$ 0.036 1.671 $\pm$ 0.064 1.492 $\pm$ 0.083 1.113 $\pm$ 0.114
@@
    UniTime 0.405 $\pm$ 0.002 37.248 $\pm$ 0.177 0.361 $\pm$ 0.002 0.166 $\pm$ 0.001 0.319 $\pm$ 0.001 0.496 $\pm$ 0.005 0.314 $\pm$ 0.002 0.352 $\pm$ 0.002 0.409 $\pm$ 0.003
    Time-LLM (ETTh1) 0.458 $\pm$ 0.002 43.014 $\pm$ 0.164 0.440 $\pm$ 0.003 0.191 $\pm$ 0.000 0.371 $\pm$ 0.000 0.499 $\pm$ 0.002 0.399 $\pm$ 0.001 0.368 $\pm$ 0.001 0.424 $\pm$ 0.001
    ChatTime-Base 0.725 $\pm$ 0.002 38.771 $\pm$ 0.351 0.837 $\pm$ 0.002 0.205 $\pm$ 0.003 1.090 $\pm$ 0.004 0.250 $\pm$ 0.004 0.365 $\pm$ 0.006 1.779 $\pm$ 0.003 2.398 $\pm$ 0.004
    ChatTime-Chat 0.781 $\pm$ 0.015 35.260 $\pm$ 0.446 0.865 $\pm$ 0.022 0.165 $\pm$ 0.002 1.217 $\pm$ 0.036 0.282 $\pm$ 0.020 0.445 $\pm$ 0.064 1.896 $\pm$ 0.001 2.536 $\pm$ 0.000